OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (–4, 89), (–3, 7), (–1, –1), (1, –1), (4, 329) a)y = 2x3 – 3x2 – 2x + 1 b)y = 1x4 – 2x3 – 3x2 + 2x + 1 c)y = x4 – 2x3 + 3x2 + 2x – 1 d)y = x4 + 2x3 – 3x2 – 2x + 1

OpenStudy (amistre64):

start narrowing the options by using the most convenient points to chk with

OpenStudy (amistre64):

the only other option would be to build the poly and thats not something you wanna do by hand with that many points

OpenStudy (anonymous):

1. What polynomial has a graph that passes through the given points? (–4, 89), (–3, 7), (–1, –1), (1, –1), (4, 329) (1 point) (0 pts) y = 2x3 – 3x2 – 2x + 1 (0 pts) y = 1x4 – 2x3 – 3x2 + 2x + 1 (0 pts) y = x4 – 2x3 + 3x2 + 2x – 1 (1 pt) y = x4 + 2x3 – 3x2 – 2x + 1 1 /1 point 2. The table shows the amount of milk that Wisconsin dairy farms produced from 1937 to 1968. Which model best fits this data? Year Milk Produced (in billions of lbs.) 1937 12 1968 16 1995 20 (1 point) (0 pts) y = –0.139x + 6.57 (0 pts) y = 6.57x + 0.139 (1 pt) y = 0.139x + 6.57 (0 pts) y = 0.139x – 6.57 1 /1 point 3. Find a model for the set of values. x –5 –4 0 4 5 y 25 16 2 16 25 (1 point) (0 pts) (0 pts) (0 pts) (1 pt) 1 /1 point Use the table below to answer questions 4–6. 4. The table shows the annual consumption of cheese per person in the United States for selected years in the 20th century. Year Pounds Consumed 1908 5.027 1923 8.768 1960 12.29 1981 15.485 What cubic model best fits this data? (1 point) (0 pts) y = –0.0000541x3 + 0.00789x2 – 0.452x – 1.888 (0 pts) y = –0.0000541x3 – 0.00789x2 – 0.452x + 1.888 (0 pts) y = 0.00789x3 – 0.0000541x2 + 1.888x + 0.452 (1 pt) y = 0.0000541x3 – 0.00789x2 + 0.452x + 1.888 1 /1 point 5. If you were going to use this model to estimate the amount of cheese consumed per person in 1976, would that prediction be considered interpolation or extrapolation? (1 point) (1 pt) interpolation (0 pts) extrapolation 1 /1 point 6. What is the domain of the graph for the data set? (1 point) (0 pts) 1976 (0 pts) 5.027 ≤ domain ≤ 15.485 (1 pt) 1908 ≤ domain ≤ 1981 (0 pts) any year in the twentieth century 1 /1 point 7. Suppose a data set follows the path of a polynomial with two turning points. Which type of polynomial function is most likely to best fit the data set? (1 point) (0 pts) linear (0 pts) quadratic (1 pt) cubic (0 pts) not enough information 1 /1 point

OpenStudy (solomonzelman):

if you want to be banned, then this should be a very good way to do it.